Operation Summary for January 13, 2010 - January 15, 2010
Wednesday Jan. 13
The day shift began with Operations monitoring store 7500 (64.48E30), the stash (129.1E10), and the Antiproton Source (Pbar) stack (18.6E10), with the Linac off due to work on a RF station (LRF2), with the Booster and Main Injector (MI) conducting maintenance or studies, with the NuMI and MiniBooNE experiments conducting maintenance or studies, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, Linac personnel completed their work on LRF2 at 2:44 PM. After LRF2 was brought online, there were a few trips as the station warmed up. A Pbar power supply (M:HV102R) failed to ramp properly following a pre-vault access earlier in the shift; support personnel investigated the problem.
During the evening shift, Operations terminated store 7500 at 7:03 PM and then put the TeV into shot setup. Operations established store 7501 at 9:03 PM with an average initial luminosity (AIL) of 156.77E30. Operations reported difficulties loading protons during shot setup, which required extensive tuning. Operations reported a Pbar LCW leak at 9:13 PM somewhere in the transport line. Operators accessed and found a burst water hose, which Water Group Techs replaced. Interlock technicians resolved a coasting beam valve problem.
Thursday Jan. 14
The midnight shift began with Operations monitoring store 7501 (89.58E30) and the stash (47.7E10), with Pbar stacking (13.0E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations reported some minor downtime during the shift due to Linac low-energy RF stations tripping off occasionally. Operations terminated store 7501 around 7:56 AM and then prepared the TeV for access.
During the day shift, Tev A-E, Transfer Hall, and the collision halls were ready for controlled access by 8:44 AM. The accesses concluded by 12:00 PM. There was some downtime due to a vacuum burst at MI location 606. PBars were stacked and transferred to the Recycler as appropriate. Beam to NuMI and MiniBooNe when it was available.
During the evening shift, Tev orbit smoothing took place from 4:00 PM to 5:00 PM. There was a quench at D1 during a proton-only wet squeeze. Recovery followed. Orbit checks were complete by 10:30 PM, shot setup followed and lasted through the end of the shift. PBars were stacked and transferred to the Recycler as appropriate. Beam to NuMI and MiniBooNe when it was available.
Friday Jan. 15
The midnight shift began with Operations in shot setup, and with NuMI and MiniBooNE taking beam.
During the midnight shift, Store 7507 was colliding by 12:15AM with an average initial luminosity of 299.1E30. Enclosures C,D,E&J were made up for Switchyard power supply testing/running. PBars were stacked and transferred to the Recycler as appropriate. Beam to NuMI and MiniBooNe when it was available. Proton beam operations were terminated at 7:00 AM and Booster was prepared for controlled access. Linac safety system testing started at the very end of the shift.
The Plans for Friday and the Weekend
The plans for today are to complete the Booster access and put in a Tev store. The plan for the weekend is to stack, stash and store.
Comments and Suggestions
|last modified 02/03/2010 email Fermilab|
|Security, Privacy, Legal|